# EE-LLM: Large-Scale Training and Inference of Early-Exit Large Language Models with 3D Parallelism

### v2

@@ 4.1.1 Methodology
To tackle this challenge, we propose a simple yet effective algorithm that instructs each pipeline stage to calculate the desired gradients correctly, without any additional communication overhead between stages. To explain our method, let us first re-write the early-exit training loss defined in Eq. (1) as
${\mathcal{L}}=\sum_{i\in[K]}{\mathcal{L}}_{i},$ (2)
where $K$ represents the number of pipeline stages, and each ${\mathcal{L}}_{i}$ is itself a weighted sum of one or multiple early/final-exit losses within Stage $i$ [^2] It is easy to check that this is equivalent to the loss defined in Eq. (1). In fact, each ${\mathcal{L}}_{i}$ can be a general objective function, as long as it is defined locally in Stage $i$; our analysis in the following still holds true in this general case. Consider one data sample $\bm{x}$ for simplicity, and each loss function is calculated with $\bm{x}$, i.e. ${\mathcal{L}}_{i}={\mathcal{L}}_{i}(\bm{x})$; in addition, let $\bm{x}_{i}$ be the hidden states that Stage $i$ calculates and sends to its next stage during the forward step. Then, during the backward step, Stage $i$ receives some gradient tensor $\bm{g}_{i}$ from Stage $i+1$, defines some auxiliary loss ${\mathcal{L}}^{\mathsf{aux}}_{i}$, and performs usual backward computation for ${\mathcal{L}}^{\mathsf{aux}}_{i}$. The auxiliary losses $\{{\mathcal{L}}^{\mathsf{aux}}_{i}\}_{i\in[K]}$ and gradient tensors $\{\bm{g}_{i}\}_{i\in[K-1]}$ are defined inductively:
${\mathcal{L}}^{\mathsf{aux}}_{K}$ $\coloneqq{\mathcal{L}}_{K},\quad\text{and}$ (3a)
${\mathcal{L}}^{\mathsf{aux}}_{i}$ $\coloneqq{\mathcal{L}}_{i}+\langle\bm{g}_{i},\bm{x}_{i}\rangle,\quad\text{where}\quad\bm{g}_{i}\coloneqq\frac{\partial{\mathcal{L}}^{\mathsf{aux}}_{i+1}}{\partial\bm{x}_{i}},\quad i=K-1,K-2,\dots,2,1.$ (3b)
Intuitively, the linear term $\langle\bm{g}_{i},\bm{x}_{i}\rangle$, i.e. the sum of entrywise product between $\bm{g}_{i}$ and $\bm{x}_{i}$, summarizes information about the gradients of all losses located in later stages. Note that $\bm{g}_{i}$ is regarded by Stage $i$ as a constant tensor, and no gradient is calculated with respect to it.
A visualization of this process can be found in Figure 2. It has the same P2P communication scheme as in the case of training a standard LLM with pipeline parallelism; the only difference is how each gradient tensors $\bm{g}_{i}$ is defined locally in Stage $i+1$. In the following, we formally prove that the proposed method leads to correct gradient calculation for the training objective in Eq. (2).
Figure 2: A high-level visualization for the process of backpropagation, including forward and backward passes, for an early-exit model partitioned into four pipeline stages.
@@ 4.1.2 Rationale
Proposition 1.
Suppose that there is no tied/shared parameter across pipeline stages, and consider the auxiliary losses defined in Eq. (4.1.1). Then, for any $i\in[K]$ and any model parameter or activation tensor $\bm{z}$ in Stage $i$, it holds that
@@
Notice that for any model parameter $\bm{z}$ in Stage $i$, one has $\partial{\mathcal{L}}_{j}/\partial\bm{z}=\bm{0}$ for any $j<i$, due to the sequential structure of early-exit LLMs (or other deep neural networks). Combining this with Eq. (2) and (4) yields
@@
which implies the correctness of gradient calculation for the training objective ${\mathcal{L}}$.
Proof of Proposition 1.
The claim of Eq. (4) is obviously true for the base case $i=K$, by definition of ${\mathcal{L}}^{\mathsf{aux}}_{K}={\mathcal{L}}_{K}$. Let us prove by induction for the remaining stages. Suppose that Eq. (4) holds true for Stage $i+1$, namely
$\frac{\partial{\mathcal{L}}^{\mathsf{aux}}_{i+1}}{\partial\bm{z}}=\frac{\partial(\sum_{j=i+1}^{K}{\mathcal{L}}_{j})}{\partial\bm{z}}.$
To prove Eq. (4) for Stage $i$, first note that by definition of $\bm{g}_{i}$, we have
@@
$\frac{\partial{\mathcal{L}}^{\mathsf{aux}}_{i}}{\partial\bm{z}}$ $=\frac{\partial({\mathcal{L}}_{i}+\langle\bm{g}_{i},\bm{x}_{i}\rangle)}{\partial\bm{z}}=\frac{\partial{\mathcal{L}}_{i}}{\partial\bm{z}}+\frac{\partial\langle\bm{g}_{i},\bm{x}_{i}\rangle}{\partial\bm{x}_{i}}\frac{\partial\bm{x}_{i}}{\partial\bm{z}}=\frac{\partial{\mathcal{L}}_{i}}{\partial\bm{z}}+\bm{g}_{i}\frac{\partial\bm{x}_{i}}{\partial\bm{z}}$
$=\frac{\partial{\mathcal{L}}_{i}}{\partial\bm{z}}+\frac{\partial(\sum_{j=i+1}^{K}{\mathcal{L}}_{j})}{\partial\bm{x}_{i}}\frac{\partial\bm{x}_{i}}{\partial\bm{z}}=\frac{\partial{\mathcal{L}}_{i}}{\partial\bm{z}}+\frac{\partial(\sum_{j=i+1}^{K}{\mathcal{L}}_{j})}{\partial\bm{z}}=\frac{\partial(\sum_{j=i}^{K}{\mathcal{L}}_{j})}{\partial\bm{z}}.$
@@
Let us move on to relax the assumption. In the broader scenario with tied/shared model parameters (e.g. word embedding matrices [54]) across pipeline stages, gradient calculation via backpropagation is equivalent to the following two-step procedure (as is implemented in Megatron-LM): (1) compute gradients as if all parameters are untied, then (2) sum up and synchronize gradients for tied parameters via all-reduce operations. Hence our proposed auxiliary-loss approach, when applied to the first part of the above two-step procedure, is still valid.
4.1.3 Choices of pipeline schedules
Our previous analysis shows how to modify the forward and backward steps for each microbatch, in order to execute backpropagation of the early-exit training loss through pipeline stages. In principle, this lightweight modification can be applied to general pipeline schedules of forward and backward steps for multiple microbatches. The classical 1F1B (one-forward-one-backward) schedule, also called PipeDream-Flush [47, 20, 49], achieves a good balance of algorithmic simplicity, training time, memory usage and communication latency, in comparison to other schedules such as GPipe [28] (larger activation memory) and interleaved 1F1B [49] (higher memory and communication requirements). Therefore, we focus on the 1F1B schedule throughout this work, for concreteness; a visualization can be found in Figure 3(a).

### v3

@@ 3.1.1 Methodology
To tackle this challenge, we propose a lightweight algorithm that instructs each pipeline stage to calculate the desired gradients correctly, without any additional communication overhead between stages. To explain our method, let us first re-write the training objective defined in Eq. (1) as ${\mathcal{L}}=\sum_{i\in[K]}{\mathcal{L}}_{i}$, where $K$ represents the number of pipeline stages, and each ${\mathcal{L}}_{i}$ is itself a weighted sum of one or multiple early/final-exit losses within Stage $i$ [^1] Obviously, the objective in Eq. (1) is a special case of this formulation. In fact, our analysis in the following allows each ${\mathcal{L}}_{i}$ to be a general objective function defined locally in Stage $i$. Consider one data sample $\bm{x}$ for simplicity, and each loss function is calculated with $\bm{x}$, i.e. ${\mathcal{L}}_{i}={\mathcal{L}}_{i}(\bm{x})$; in addition, let $\bm{x}_{i}$ be the hidden states that Stage $i$ calculates and sends to its next stage during the forward step. Then, during the backward step, Stage $i$ receives some gradient tensor $\bm{g}_{i}$ from Stage $i+1$, defines some auxiliary loss ${\mathcal{L}}^{\mathsf{aux}}_{i}$, and performs usual backward computation for ${\mathcal{L}}^{\mathsf{aux}}_{i}$. The auxiliary losses $\{{\mathcal{L}}^{\mathsf{aux}}_{i}\}_{i\in[K]}$ and gradient tensors $\{\bm{g}_{i}\}_{i\in[K-1]}$ are defined inductively:
${\mathcal{L}}^{\mathsf{aux}}_{K}$ $\coloneqq{\mathcal{L}}_{K};\quad\text{for }i=K-1,K-2,\dots,1,$ (2a)
${\mathcal{L}}^{\mathsf{aux}}_{i}$ $\coloneqq{\mathcal{L}}_{i}+\langle\bm{g}_{i},\bm{x}_{i}\rangle,\quad\text{where}\quad\bm{g}_{i}\coloneqq\frac{\partial{\mathcal{L}}^{\mathsf{aux}}_{i+1}}{\partial\bm{x}_{i}}.$ (2b)
Intuitively, the linear term $\langle\bm{g}_{i},\bm{x}_{i}\rangle$, i.e. the sum of entrywise product between $\bm{g}_{i}$ and $\bm{x}_{i}$, summarizes information about the gradients of all losses located in later stages. Note that $\bm{g}_{i}$ is regarded by Stage $i$ as a constant tensor, and no gradient is calculated with respect to it. A visualization of this process can be found in Figure 2. It has the same P2P communication scheme as in the case of training a standard LLM with pipeline parallelism; the only difference is how each gradient tensors $\bm{g}_{i}$ is defined locally in Stage $i+1$. In the following, we prove that the proposed method leads to correct gradient calculation for the training objective ${\mathcal{L}}$.
Figure 2: The backpropagation process for an early-exit model partitioned into four pipeline stages.
@@ 3.1.2 Rationale
Proposition 3.1.
Suppose that there is no tied parameter across pipeline stages, and consider the auxiliary losses defined in Eq. (2). Then, for any $i\in[K]$ and any model parameter or activation tensor $\bm{z}$ in Stage $i$, it holds that
@@
Notice that for any model parameter $\bm{z}$ in Stage $i$, one has $\partial{\mathcal{L}}_{j}/\partial\bm{z}=\bm{0}$ for any $j<i$, due to the sequential structure of early-exit LLMs (or other deep neural networks). Combining this with ${\mathcal{L}}=\sum_{i\in[K]}{\mathcal{L}}_{i}$ and Eq. (3) yields
@@
implying correctness of gradient calculation for ${\mathcal{L}}$.
Proof of Proposition 3.1.
The claim of Eq. (3) is obviously true for the base case $i=K$, by definition of ${\mathcal{L}}^{\mathsf{aux}}_{K}={\mathcal{L}}_{K}$. Let us prove by induction for the remaining stages. Suppose that Eq. (3) holds true for Stage $i+1$, namely ${\partial{\mathcal{L}}^{\mathsf{aux}}_{i+1}}/{\partial\bm{z}}={\partial(\sum_{j=i+1}^{K}{\mathcal{L}}_{j})}/{\partial\bm{z}}$. To prove Eq. (3) for Stage $i$, first note that by definition of $\bm{g}_{i}$, we have
@@
$\frac{\partial{\mathcal{L}}^{\mathsf{aux}}_{i}}{\partial\bm{z}}$ $=\frac{\partial({\mathcal{L}}_{i}+\langle\bm{g}_{i},\bm{x}_{i}\rangle)}{\partial\bm{z}}=\frac{\partial{\mathcal{L}}_{i}}{\partial\bm{z}}+\frac{\partial\langle\bm{g}_{i},\bm{x}_{i}\rangle}{\partial\bm{x}_{i}}\frac{\partial\bm{x}_{i}}{\partial\bm{z}}$
$=\frac{\partial{\mathcal{L}}_{i}}{\partial\bm{z}}+\bm{g}_{i}\frac{\partial\bm{x}_{i}}{\partial\bm{z}}=\frac{\partial{\mathcal{L}}_{i}}{\partial\bm{z}}+\frac{\partial(\sum_{j=i+1}^{K}{\mathcal{L}}_{j})}{\partial\bm{x}_{i}}\frac{\partial\bm{x}_{i}}{\partial\bm{z}}$
$=\frac{\partial{\mathcal{L}}_{i}}{\partial\bm{z}}+\frac{\partial(\sum_{j=i+1}^{K}{\mathcal{L}}_{j})}{\partial\bm{z}}=\frac{\partial(\sum_{j=i}^{K}{\mathcal{L}}_{j})}{\partial\bm{z}}.$
@@
Let us move on to relax the assumption. In the broader scenario with tied model parameters, e.g. word embedding matrices (Press & Wolf, 2017), across pipeline stages, gradient calculation via backpropagation is equivalent to the following two-step procedure: (1) compute gradients as if all parameters are untied, then (2) sum up and synchronize gradients for tied parameters via all-reduce operations. Hence our proposed auxiliary-loss approach, when applied to the first part of this two-step procedure, is still valid.
Figure 3: One iteration of the 1F1B pipeline schedule, in a setting with $P=4$ pipeline stages and $M=6$ microbatches per batch. At the top of this figure, “Backbone forward/backward” stands for computation of Transformer layers on the backbone, while “Exit forward/backward” stands for computation of early-exit or final-exit layers. The number in each block denotes the index of the corresponding microbatch. Critical paths are marked by dashed red lines. From Figure (a) to (b), additional “Exit forward/backward” blocks are added, due to the introduction of early exits to middle stages. From Figure (b) to (c), the order of computation is slightly adjusted for the purpose of reducing memory usage. For clarity, we ignore computation related to the input embedding layer, and P2P communication latency between pipeline stages.
3.1.3 Integration with the 1F1B schedule
We have shown how to modify the forward and backward steps for each microbatch, in order to execute backpropagation of the early-exit training loss through pipeline stages. In principle, this lightweight modification can be integrated with general pipeline schedules of forward and backward steps for multiple microbatches. The classical 1F1B (one-forward-one-backward) schedule, also called PipeDream-Flush (Narayanan et al., 2019; Fan et al., 2021; Narayanan et al., 2021b), achieves a good balance of algorithmic simplicity, training time, memory usage and communication latency, in comparison to other schedules such as GPipe (Huang et al., 2019) (larger activation memory) or interleaved 1F1B (Narayanan et al., 2021b) (higher memory and communication requirements). Hence for concreteness, we focus on the 1F1B schedule throughout this work. Within one iteration of this schedule, each stage goes through a warm-up phase (forward steps of the beginning microbatches), a steady 1F1B phase, and a cool-down phase (backward steps of the final microbatches). A visualization can be found in Figure 3(a).
@@ 3.2 Training efficiency
At first glance, one might expect that adding early exits to a standard LLM will incur a training overhead, in terms of time and memory, that is (at least) proportional to the number of additional model parameters. Fortunately, this is not the case for training with pipeline parallelism. We discover that computation related to the additional early exits can leverage idle resources in the original 1F1B schedule, thus causing negligible overhead to training time and zero overhead to peak GPU memory under mild conditions.
To explain this, let us first identify three types of idle resources in the original 1F1B schedule.
  • Explicit bubbles, i.e. the light gray areas in Figure 3(a), during which GPUs are idle.
  • Implicit bubbles, i.e. the dark gray areas in Figure 3(a), caused by load imbalance across pipeline stages. In particular, the last pipeline stage requires additional computation for the output layer.
  • Idle memory, caused by unbalanced memory usage: earlier stages have to save the intermediate activations for more microbatches, and the first/last stage need to store an extra input/output layer, plus the corresponding gradients and optimizer states. As a result, the first stage is typically the bottleneck of peak memory usage.
Now, suppose that we choose $k$ middle stages and add one minimalistic early-exit layer to each of them. As shown in Figure 3(b), by utilizing implicit bubbles, computation of $k$ early exits will increase the training time per iteration by only $k\times(f_{\mathsf{EE}}+b_{\mathsf{EE}})$, where $f_{\mathsf{EE}}$ (resp. $b_{\mathsf{EE}}$) represent the time for one forward (resp. backward) pass of one microbatch for one early-exit layer. Moreover, memory usage by model parameters for each middle stage will not exceed that of the first or last stage. Deferring forward computation of early exits to backward steps, as shown in Figure 3(c), further decreases the memory overhead by early-exit logits in Stage $i\in[P]$ from $s\times b\times V\times(P-i+1)$ to $s\times b\times V$, where $s$ is the sequence length, $b$ is the microbatch size, $V$ is the vocabulary size, and $P-i+1$ is the number of in-flight microbatches (Korthikanti et al., 2022) for Stage $i$. Putting these together, the peak memory usage across stages remains unchanged, as long as $s\times b\times V$ is less than the activation memory of all backbone Transformer layers within one stage for one microbatch.
Due to limited space, we defer detailed and formal analysis of training efficiency in broader settings to Appendix A.
3.3 Advanced features
EE-LLM incorporates some advanced features that can potentially improve the training process. One of them is the option of changing early-exit loss weights during training, just like learning rate or other hyperparameters. Another one is a novel method of improving resource utilization by filling explicit bubbles with partial forward and backward computation for additional microbatches, which is visualized in Figure 4. One can prove formally that such additional computation provides useful gradient information for optimizing the early-exit training objective, without increasing the training time per iteration. More details of both features can be found in Appendix C.
Figure 4: The proposed method of filling bubbles with additional microbatches. In this example, P1 and P2 go through the forward and backward passes for the first few stages, while P3 and P4 go through the full forward pass, followed by the backward pass for the last few stages.
4 Inference
This section first explains the major challenge faced by early-exit LLM inference for autoregressive generation, and a few recent attempts to resolve it. Then, we introduce a novel solution based on a new type of pipeline parallelism. Throughout this section, we focus on the latency-oriented setting, with a batch size of 1 during sequence generation.
Main challenge: KV caching.
Accelerating inference with early exiting is, in principle, orthogonal to and compatible with many other common techniques of acceleration, such as kernel fusion, FlashAttention (Dao et al., 2022), quantization (Yao et al., 2022; Xiao et al., 2023), among others (Kim et al., 2023a). For autoregressive generation, one exception is KV caching, i.e. saving keys and values of all attention layers for previously generated tokens, which reduces redundant computation at the cost of higher memory usage. This is contradictory to vanilla early-exit inference: if the current token is generated via early exiting, then its KV caches in later layers are missing, which hinders the generation of future tokens that go beyond the exiting layer of the current token. This challenge has been well recognized in the literature, and several approaches have been recently proposed to resolve it, including state propagation (Elbayad et al., 2020; Li et al., 2021; Schuster et al., 2022), SkipDecode (Corro et al., 2023), and synchronized parallel decoding (Bae et al., 2023; Tang et al., 2023). See Appendix F for more discussion on these methods.
A variant of the last method, which we call KV recomputation, is implemented in EE-LLM. In this approach, we maintain a list of the most recent tokens that have missing KV caches in deep layers due to early exiting. During each forward pass, we include these early-exit tokens in the current forward pass, which allows for direct recomputation of the KV caches for these tokens and thus avoids the issue of missing KV caches. It relies on the batching effect of GPU computation for early-exit acceleration, yet might not achieve any acceleration on other hardware platforms, due to its high theoretical complexity.
Figure 5: Standard full-model inference (top) and our pipeline-based early-exit inference (bottom). Numbers in the blocks denote the tokens within one generated sequence. For simplicity of visualization, we assume here that (1) each early exit is located at the end of some pipeline stage, and (2) the latency for generating each token is the same (while in practice, generating the first token via the prefilling phase usually takes longer than generating another token during the decoding phase).
New solution: pipeline parallelism.
We propose a novel type of pipeline parallelism to tackle the aforementioned challenge. The key idea is that, in the process of inference with multiple pipeline stages, the following two processes run in parallel whenever the model decides to do early exiting for the current token at a certain exit:
  • The full-model forward pass of the current token is continued from the exiting layer, which fulfills its KV caches in all later layers.
See Figure 5 for a visualization of this approach. Although each token essentially goes through a forward pass of the full model, the computation after the exiting layer is parallelized with the computation of later tokens, which is how acceleration is achieved in this approach. It can be checked that the inference latency for generating one token at a certain exit matches exactly the time needed for the forward computation before returning an output at that exit, unless the selected exit is located in the middle of the first pipeline stage, in which case generation of the next token has to wait until the forward pass of the first stage for the current token is completed. Note that this is true not just in practice but also for the theoretical time complexity, without relying on the batching effect of GPU computation like KV recomputation does. An implementation of this pipeline-based inference method is provided in EE-LLM. One potential limitation of the proposed method is that it requires multiple devices to facilitate pipeline parallelism, although parallelism within a single GPU or other device might be possible with more advanced implementation.
5 Experiments
5.1 Training
In the following experiments, we empirically investigate the convergence of training early-exit models with EE-LLM, as well as the training efficiency of EE-LLM for early-exit LLMs up to an unprecedented scale of 30B. This scale is only limited by the hardware resources available to us, namely an 8-node cluster with 8 Nvidia A100-80GB GPUs in each node and hence 64 GPUs in total. We use a random subset of the pre-training data provided by Data-Juicer (Data-Juicer, 2023; Chen et al., 2023). Experiments for models of the same size use the same subset and order of data. In all cases, we use the Adam optimizer (Kingma & Ba, 2014) with $\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=10^{-8}$, and a cosine schedule for the learning rate, with a maximum value of $3\times 10^{-4}$.
Convergence of training losses.
We first consider a 1.3B GPT Transformer with 24 layers, add one minimalistic early-exit layer without layer normalization to the 1/4 depth and the other to the 1/2 depth, set their early-exit loss weights to 1/4 and 1/2 respectively (while the final-exit loss has a weight of 1), and tie all input and output embedding matrices. A standard LLM of the same architecture is trained using the same hyperparameters and pre-training data. We further train a 7B early-exit model with 32 layers using similar configurations, except that early-exit loss weights are set to 0.1 and 0.2, and all embedding matrices are untied. A standard 7B model is trained similarly. The batch size and sequence length are both set to 2048.
(a) 1.3B, 24 layers
(b) 7B, 32 layers
Figure 6: Convergence of early-exit/final-exit training losses. Each curve is annotated with the index of the Transformer layer that the corresponding exit is connected to.
Figure 6 shows the convergence of early and final-exit training losses, i.e. negative log-likelihood of next-token prediction, for both standard and early-exit LLMs. All loss curves decay at a similar pace, and unsurprisingly, the early-exit losses are slightly higher than the final-exit loss for each model. Interestingly, the final-exit loss curve of each early-exit model is close to (or even slightly below) that of the standard model, suggesting that optimizing for early-exit losses might not hurt the full-model output in our setting. We also observe more spikes in the loss curves for the 7B models than for the 1.3B models, possibly because (1) we choose to untie the embedding matrices and use smaller early-exit loss weights, both of which incur weaker regularization for the training process; and (2) layer normalization, which is known to stabilize the training of LLMs, is not included in the minimalistic early-exit layers of our 7B model. Empirical results for the convergence of training losses under more general early-exit configurations can be found in Appendix B.3.
(a) 1.3B
(b) 7B
(c) 13B
(d) 30B
Figure 7: Training time per iteration and peak GPU memory vs. the number of added early exits. Note that wall-clock time can be impacted by other workloads on the same GPU cluster when our experiments were conducted, which inevitably causes perturbations to our numerical results.
Training efficiency.
Starting with a standard GPT Transformer of size ranging from 1.3B to 30B, we increase the number of early exits from 0 to 3. Minimalistic early exits are added one by one to specific locations in the following order: (1) to the 1/4 depth; (2) to the 1/2 depth; (3) to the hidden states right before the first Transformer layer, which is always located in the first pipeline stage. Based on the performance optimizations proposed in Appendix A, when an early exit is inserted into the middle of two layers located on two consecutive pipeline stages, we always add it to the beginning of the latter stage. We set the global batch size to 2048, microbatch size to 2 (for 1.3B and 7B models) or 1 (for 13B and 30B models), and sequence length to 2048. The data parallelism degree is fixed at 4, while tensor (TP) and pipeline (PP) parallelism degrees take various values.
The empirical results illustrated in Figure 7 matches our analytical study in Section 3.2. In particular, training time increases with the number of added early exits, but at a slower rate when pipeline parallelism is enabled, thanks to the proposed utilization of implicit bubbles. Without pipeline parallelism, peak GPU memory increases with the number of early exits; on the other hand, with the pipeline parallelism degree set to 4, peak memory remains unchanged as early exits are added to middle pipeline stages, and only increases when the last early exit is added to the first stage.
5.2 Inference
In the experiments below, we verify the downstream performance of the pipeline-based approach proposed in Section 4, with the number of pipeline stages set to 4. A server with 4 Nvidia A100-40GB GPUs is used for inference. We use the 1.3B and 7B early-exit LLMs from the previous experiment on convergence of training losses in Section 5.1, which have been pre-trained from scratch using 300B and 150B tokens respectively, without further fine-tuning or alignment. Since designing the decoding method or exit condition is not our focus, we simply adopt greedy decoding and a basic confidence-based exit condition: at each early exit, if the maximum probability of next-token prediction is above a pre-defined threshold, then the model chooses the most likely token and moves on to generate the next one. The trade-off between output quality and speed is controlled by this threshold. When the threshold is set to 1, early-exit layers are disabled, so that the time complexity of a full-model forward pass matches that of a standard LLM [^2] We observed empirically that the actual wall-clock latency between these two cases might differ by up to 2%, which is negligible compared to the early-exit speedup.. Inference latency in this case is used as the baseline for calculating relative speedup. For benchmarking, we evaluate our models with HELM (Liang et al., 2023) in six tasks: BoolQ (Clark et al., 2019), TruthfulQA (Lin et al., 2022), NaturalQuestions open-book and closed-book (Kwiatkowski et al., 2019), XSUM (Narayan et al., 2018), and CNN/DailyMail. The first four tasks are question-answering tasks, while the last two are summarization tasks. We use EM (ratio of exact match with the correct answers) as the metric for BoolQ and TruthfulQA, F1 for NaturalQuestions open-book/closed-book, and ROUGE-L (a measure of similarity to the reference summaries) for XSUM and CNN/DailyMail.
(a) 1.3B
(b) 7B
Figure 8: Evaluation scores and relative speedup with our early-exit models. For each curve, speedup increases as the confidence threshold decreases from left to right. The right column also includes horizontal dotted lines that represent the performance of a 7B standard LLM trained under the same configurations as those for our 7B early-exit model.
Figure 8 demonstrates the scores and speedup for pipeline-based inference with varying confidence thresholds. Encouragingly, in many cases, inference with early exiting achieves 2×2$\times$2 × or higher speedup compared to full-model inference, with comparable or even better evaluation scores. Table 3 provides example texts generated by our 7B early-exit model, showing that speedup can be achieved with minor or no impact on the resulting sequence. Table 4 further demonstrates the confidence at each exit for each token, confirming the existence of easy tokens that can be predicted correctly with high confidence at early exits. See also Appendix B.1 for an empirical comparison between the pipeline-based method and KV recomputation, which are shown to perform similarly in our setting. It is worth noting that performance of early-exit inference might be improved in many ways, e.g. using early exits of different structures, other decoding mechanisms and exit conditions, or simply more extensive and sufficient pre-training. We also conjecture that early exiting can, in principle, bring higher speedup for larger LLMs, as expressivity of early exits grows with the model capacity, which allows them to make confident and accurate predictions for a larger proportion of tokens.
6 Conclusions
We have introduced EE-LLM, a system for large-scale training and inference of early-exit LLMs with 3D parallelism. For training, we have presented how to execute backpropagation of the early-exit training objective across pipeline stages, performance optimizations for minimizing the computational overhead compared to training standard LLMs, and advanced features for fine-grained control and optimization of the training process. For inference, we have introduced our design and implementation of two inference methods, one based on KV recomputation and the other based on a new type of pipeline parallelism, both of which are compatible with KV caching for autoregressive generation. Along the way, we have discovered some interesting chemical reactions between early exiting and pipeline parallelism, which is likely because it is both along the depth dimension that an early-exit model executes forward computation selectively and gets partitioned into pipeline stages. We hope that EE-LLM will be a helpful tool for future study and applications of early-exit LLMs at scales as large as those of standard LLMs. It is worth noting that many ideas presented in this work can be extended to more general neural network architectures, deep learning frameworks, hardware platforms, or other broader settings. It would be exciting to see further developments in these aspects.
Acknowledgements
We would like to thank the reviewers for their constructive and valuable feedback that helps improve this work.
Impact Statement
This paper presents work whose goal is to advance techniques for training and inference of early-exit large language models. Therefore, it bears potential societal consequences similar to those of prior literature on the technical aspects of LLM training and inference, e.g. misuse of the resulting pre-trained LLMs. Other than that, we see nothing that we feel must be specifically highlighted here.
Appendix A Training efficiency